# envname-P envname#1

### v1

@@ Abstract.
Supply chain attacks threaten open-source software ecosystems. This paper proposes a formal framework for quantifying trust in third-party software dependencies that is both formally checkable – formalized in satisfiability modulo theories (SMT) – while at the same time incorporating human factors, like the number of downloads, authors, and other metadata that are commonly used to identify trustworthy software in practice. We use data from both software analysis tools and metadata to build a first-order relational model of software dependencies; to obtain an overall “trust cost” combining these factors, we propose a formalization based on the minimum trust problem which asks for the minimum cost of a set of assumptions which can be used to prove that the code is safe. We implement these ideas in Cargo Sherlock, targeted for Rust libraries (crates), incorporating a list of candidate assumptions motivated by quantifiable trust metrics identified in prior work. Our evaluation shows that Cargo Sherlock can be used to identify synthetically generated supply chain attacks and known incidents involving typosquatted and poorly AI-maintained crates, and that its performance scales to Rust crates with many dependencies.
@@ 1. Introduction
As software supply chain ecosystems grow in size and complexity, automated techniques are needed to prevent supply chain attacks (ohm2020backstabber, ; martinez2021software, ). Recent and notable examples of such attacks include the SolarWinds attack (martinez2021software, ); the XZ Utils backdoor (openwallOsssecurityBackdoor, ; helpnetsecurityBewareBackdoor, ); and for Rust, which is the focus of our implementation, the malicious libraries rustdecimal (rustsecRUSTSEC20220042Rustdecimal, ) and faster_log (fasterlog-asyncprintln, ). These attacks occur at the package granularity, where untrustworthy libraries or software dependencies are added to trustworthy projects, or safe libraries are hijacked by malicious developers. Many existing techniques can be used to help identify such attacks. Some of these are formally checkable in the sense that they are based on a well-defined formal model of software: e.g., static analysis tools, fuzzers, type systems, and verification tools. For such tools, analysis results correspond to concrete syntactic or semantic source code properties or program traces, and can therefore be understood in an appropriate formal system (moller2012static, ). However, these tools are often limited to specific semantic or syntactic properties of programs.
@@
On the other hand, industry developers and security engineers typically rely on a rather different set of criteria in practice: they use a combination of trust in software and auditing to ensure that third-party dependencies are safe from software supply chain attacks (hamer2025trusting, ; cargo-vet, ; cargo-scan-paper, ; jfrogClosingSupply, ; crev, ). For example, developers at Google may be expected to trust software that was developed by the company or one of its industry partners; at a smaller scale, an independent developer may be more inclined to trust software that is widely accepted by the community (e.g., has a large number of downloads and stars on GitHub). Trust and auditing are the foundation of techniques like software bill of materials (SBOM) (muiri2019framing, ; hendrick2022state, ; xia2023empirical, ; zahan2023software, ), which require human-curated or manually generated lists of software components and their purposes. However, trust and auditing (as currently practiced) often come with no formal guarantees; there is no formal model over which we can understand the meaning of audits and trust relationships, and how they compose.
@@
Research Problem and Goals. Is it possible to quantify trust in a way that is formally checkable (in some formal system) yet also accounts for human factors that many developers rely on in practice? Our specific focus is on quantifying trust in a first-order logic framework; roughly speaking, the more assumptions that must be made in a software dependency and taken “on faith”, the greater the trust cost that we should need to pay in order to believe it functions correctly. It should be the case that higher trust costs correspond to more risky code. We identify the following specific goals for such a trust cost:
@@
  • Trust should be auditable: it should be possible for a human to read and understand the assumptions made by a trust analysis, and revise these assumptions if necessary.
  • Trust should be compositional; we should be able to compute trust costs for different portions of the package ecosystem independently, and combine these trust costs.
@@
  • Finally, trust should be configurable; the user should be able to add and modify assumptions, e.g., adding additional tool outputs or human factors as parameters.
@@
We note that traditional software vulnerability analysis and other machine learning (ML) approaches (krsul1998software, ; ghaffarian2017software, ) are a non-solution to the above. Such ML techniques would meet our criteria for compositionality and configurability; but they would fail the second and third criteria, auditability and specific assumptions that can later be vetted by human developers and human users.
@@
Cargo Sherlock: An SMT-Based Approach. To address these goals, we propose Cargo Sherlock, a design and implementation for quantifying supply chain trust based on a formalization in satisfiability modulo theories (SMT). Cargo Sherlock is designed to be used directly by end users: it can be used to quantify trust when selecting dependencies to add to a project, and its features are motivated by this intended use case; see Section˜2 and Figure˜1 for an overview.
@@
We begin with an SMT-based formalization which is based on a particular formalization of trust that we call the minimum trust problem (Section˜3). The minimum trust problem formalization assumes that we have a list of assumptions that the user might make about the code (such as, if the code has a large number of downloads, then it is safe, or if it is audited by our organization, then it is safe); each assumption has an associated cost, which can be thought of as the cost to make that assumption. These assumptions are motivated by prior empirical research; they are not intended to be “true facts”, but rather, they are real thought processes that users employ to conclude that dependencies are safe, and are configurable by users. The minimum trust problem then is to determine the minimum cost of a set of assumptions which can be used to prove (in the sense of propositional logic) that the code is safe. We argue that this model can meet the goals we outlined above; for example, it produces trust costs that are specific to a particular set of assumptions, can be configured to meet specific users’ needs, and are auditable by human developers. We consider two algorithms for solving this problem, based on a reduction to SMT. We also show that the problem is NP-hard in general. This formalization can be compared to existing work in database theory, for example, the semiring framework for database provenance (green2017semiring, ) or assigning weights to logical conclusions in Datalog (bistarelli2008weighted, ).
@@
For the design and implementation of the tool itself (Section˜4), Cargo Sherlock gathers software supply chain metadata and the results of automated program analysis tools to build a first-order relational structure of the software supply chain – that is, of all the dependencies of a given package. It then instantiates the minimum trust problem above on a set of possible assumptions on the variables involved in this first-order structure. The model also supports “parameterized assumptions”, where the assumption cost depends in a continuous way on a particular piece of metadata, and “negative assumptions”, which can be used to incorporate information about software components that are distrusted, such as known vulnerabilities or bugs found by the program analysis tools.
@@
Our evaluation (Section˜5) considers four questions: can Cargo Sherlock be used to prevent synthetically generated supply chain attacks? Would it identify risk in Rust dependencies (crates) that have been involved in known Rust supply chain incidents? How does it fare on known vulnerable crates? And lastly, does the performance of Cargo Sherlock scale to Rust crates with many dependencies? We find that synthetically generated typosquatting attacks generally result in much higher trust costs. We also find that of the two algorithms we propose, our more optimized algorithm provides timely results in crates with a large number of dependencies.
@@
One way to put these results into context is to think about a particular user model of how we expect users to interact with our code. For example, due to the higher trust costs, if using our tool in the prescribed model, a user would catch any of the typosquatting attacks considered in our evaluation by identifying and omitting the risky dependency prior to its inclusion.
@@
  • A formal model of the supply chain trust problem and two algorithms for its solution using the SMT theory of linear arithmetic (Section˜3);
  • A design and architecture for a usable tool based on these ideas, including concrete input data about Rust code and a collection of built-in candidate assumptions motivated by human factors and available program analyses (Section˜4);
  • An evaluation which studies the efficacy of our tool for identifying synthetic and real-world supply chain attacks, and its scalability on Rust libraries with many dependencies (Section˜5).
We conclude with related work (Section˜6). Cargo Sherlock is open source and released under the MIT License and is publicly available on GitHub. [^1] https://github.com/DavisPL/cargo-sherlock
@@ 2. Motivating Overview
Figure 1. Cargo Sherlock example output on the anyhow-1.0.97 crate.:
This section presents a motivating overview of Cargo Sherlock from a user-facing perspective; the reader who is more interested in formal details may skip to Section˜3. We consider an example of Ferris, a Rust developer who wishes to install a Rust crate called aho_corasick (implementing the Aho-Corasick algorithm), and explore how Cargo Sherlock’s approach might help Ferris attain the goals laid out in the introduction.
@@ 2.1. Automating trust costs
After coming across the crate aho_corasick from a Google search, Ferris has a choice: do they trust the crate blindly, or do they try to determine first if the crate is safe to install in their system? If Ferris is a typical developer, they might employ a common middle ground by relying on various heuristics to argue that the crate seems reasonably safe, e.g., that the crate has many downloads, or that it is developed by someone who is trustworthy. At its core, our tool supports Ferris in this effort by attempting to automate Ferris’ argumentation. The premises of Ferris’ arguments are based on assumptions about the supply chain ecosystem (e.g., “the more downloads a crate has, the safer it is”). Cargo Sherlock includes a default list of such assumptions motivated by prior research (see Section˜4.3). Ferris may find some assumptions more reasonable than others; we mark each assumption with a cost to reflect this reasonableness. If Ferris runs Cargo Sherlock on aho_corasick, it will present them with aho_corasick’s trust cost: the total cost of the argument whose premises have the lowest total cost to assume (informally, the strongest argument), along with the assumptions that were used to determine this trust cost. After using Cargo Sherlock, Ferris determines that aho_corasick is marked as trusted, and that it was marked SAFE due to having a high number of downloads. See Figure˜1 for an example output along these lines.
@@ 2.2. Trust should be specific
Now suppose that Ferris becomes concerned about trusting crates with a high number of downloads. Concerned, Ferris hires an expert auditor, Crabby, who combs through every line of source code of aho_corasick. Since Ferris trusts Crabby’s expertise, Ferris assigns a very low cost to the assumption that “if the crate is audited by Crabby, it is safe” – lower than any other available assumption. Now suppose that Crabby comes back after the audit and reports that the crate is completely safe to use. In this scenario, it would seem that the previous assumptions used to mark the crate trusted – such as its high number of downloads or the author of the crate – are no longer relevant to Ferris. Since Cargo Sherlock only reports the minimum trust cost for a crate with the associated assumptions, it does not present other (potentially unhelpful) assumptions which might be used to trust the crate. In other words, Ferris is only presented with a specific reason to trust the crate, and not other potential reasons. We provide a default list of assumptions and costs that are suitable for the common use cases, motivated by prior work and empirical studies; the list of assumptions we support is given in Section˜4.
@@ 2.3. Trust should be auditable and configurable
Suppose that Ferris later finds out that Crabby is untrustworthy. At this point, Ferris can return to their assumption list and correct the problem by increasing the cost of the assumption “if the crate is audited by Crabby, it is safe”. At this point, Cargo Sherlock will update its result: the trust cost for aho_corasick will increase, and the associated minimum cost set of assumptions will update. Similarly, Ferris may want to update the list of assumptions they make based on new information (like which specific authors and auditors are considered trusted).
@@ 2.4. Trust should be composable
After installing the aho_corasick crate, Ferris discovers that it depends on the memchr crate. Concerned, Ferris feels that the trust cost for aho_corasick should be dependent on the trust cost of its dependencies in some cases. One such case is if Crabby did not audit aho_corasick, as then Ferris may be concerned about the number of packages being incorporated. One way to model this is to include an assumption such as the following, which composes information about crates and their dependencies: “if a crate’s dependencies are all safe to use, and the crate itself does not have any code flagged by a static analysis tool, then the crate is safe.” Our tool uses a similar assumption to this one to allow arguments which compose trust costs on all of a crate’s dependencies to obtain a trust cost for the crate itself.
@@ 2.5. Incorporating known risks
Finally, Ferris – while browsing online – discovers the RustSec Advisory Database (rustsecAdvisoriesx203A, ) (which lists vulnerabilities for different Rust crates). Ferris wants to check whether aho_corasick is present in the database. Unfortunately, none of the aspects so far accommodates negative factors which would decrease trust in a crate; intuitively, this is because we have only incorporated components that are about trustworthiness. We thus extend the tool with a distrust model which quantifies known risks; we combine the results of the trust and distrust models to obtain a final label for the crate. This is described in Section˜4.3.
@@ 3. Minimum Trust Problem
In this section, we present the formal model behind Cargo Sherlock, which aims to satisfy the requirements that we laid out in Sections˜1 and 2. We define the minimum trust problem formally, which asks the following: given a set of candidate assumptions and a conclusion, find the set of assumptions with minimum cost that entails the conclusion. We begin with an example, followed by two algorithms for the minimum trust problem: first a naive algorithm (Algorithm˜1) and then one based on Horn clause inference (Algorithm˜2) which forms the core theoretical basis for our trust solver implementation. Finally, we show that the complexity of the minimum trust problem for Horn clauses is NP-complete, justifying such SMT-based approaches.

### v2

# Cargo Sherlock: An SMT-Based Checker for Software Trust Costs

@@ Abstract.
Supply chain attacks threaten open-source software ecosystems. This paper proposes a framework for quantifying trust in third-party software dependencies that is formally checkable – formalized in satisfiability modulo theories (SMT) – while at the same time incorporating human factors, like the number of downloads, authors, and other metadata that are commonly used to identify trustworthy software in practice. We use data from software analysis tools and metadata to build a first-order relational model of software dependencies; to obtain an overall “trust cost” combining these factors, we propose a formalization based on the minimum trust problem, which asks for the minimum cost of a set of assumptions that can be used to conclude that the code is safe. We implement these ideas in Cargo Sherlock, targeted for Rust libraries (crates), incorporating a list of candidate assumptions motivated by quantifiable trust metrics identified in prior work. Our evaluation shows that Cargo Sherlock can identify synthetically generated supply chain attacks and known incidents involving typosquatted and poorly AI-maintained crates. Its performance scales to Rust crates with many dependencies.
Formal Methods, Trust in Software, Human Factors in Security, Software Supply Chain
journalyear: 2026copyright: ccconference: IEEE/ACM 14th International Conference on Formal Methods in Software Engineering; April 12–13, 2026; Rio de Janeiro, Brazilbooktitle: IEEE/ACM 14th International Conference on Formal Methods in Software Engineering (FormaliSE ’26), April 12–13, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793656.3793687isbn: 979-8-4007-2478-7/2026/04ccs: Software and its engineering Formal methodsccs: Security and privacy Software and application securityccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
As software ecosystems grow in size and complexity, automated techniques are needed to prevent supply chain attacks (Ohm et al., 2020; Martínez and Durán, 2021). Recent and notable examples of such attacks include the SolarWinds attack (Martínez and Durán, 2021); the XZ Utils backdoor (Freund, 2024; Zorz, 2024); and for Rust, which is the focus of our implementation, the malicious libraries rustdecimal (Rust Secure Code Working Group, 2024b) and faster_log (Pearce, 2025). These attacks occur at the package granularity, where untrustworthy libraries or software dependencies are added to trustworthy projects, or safe libraries are hijacked by malicious developers. Many existing techniques can be used to help identify such attacks. Some of these are formally checkable in the sense that they are based on a well-defined formal model of software: e.g., static analysis tools, fuzzers, type systems, and verification tools. For such tools, analysis results correspond to concrete syntactic or semantic source code properties or program traces, and can therefore be understood in an appropriate formal system (Møller and Schwartzbach, 2012). However, these tools are often limited to specific semantic or syntactic properties of programs.
@@
On the other hand, industry developers and security engineers typically rely on a rather different set of criteria in practice: they use a combination of trust in software and auditing to ensure that third-party dependencies are safe from software supply chain attacks (Hamer et al., 2025; Mozilla, 2024; Zoghbi et al., 2026; Chin, 2023; development team, 2024b). For example, developers at a large company may be likely to trust software that was developed by that company or one of its industry partners; at a smaller scale, an independent developer may be more inclined to trust software that is widely accepted by the community (e.g., has a large number of downloads and stars on GitHub). Trust and auditing are the foundation of techniques like software bill of materials (SBOM) (Muirí, 2019; Hendrick and Zemlin, 2022; Xia et al., 2023; Zahan et al., 2023), which require human-curated or manually generated lists of software components and their purposes. However, trust and auditing (as currently practiced) often come with no formal guarantees; there is no formal model over which we can understand the meaning of audits and trust relationships, and how they compose.
Research goals.
@@
Is it possible to quantify trust in a way that is formally checkable (in some formal system) yet also accounts for human factors that many developers rely on in practice? Our specific focus is to quantify trust in propositional and first-order logic; roughly speaking, the more assumptions that must be taken for a software dependency “on faith”, the greater the trust cost that we should need to pay in order to believe it functions correctly. It should be the case that higher trust costs correspond to more risky code. We identify the following specific goals for such a trust cost:
@@
  • Trust should be auditable: it should be possible for a human to read and understand the assumptions made by a trust analysis.
@@
  • Trust should be configurable: the user should be able to add and revise assumptions, e.g., by adding additional tool outputs or human factors as parameters.
  • Finally, trust should be composable: we should be able to compare and combine trust costs corresponding to different portions of the package ecosystem.
@@
We note that traditional software vulnerability analysis and other machine learning (ML) approaches (Krsul, 1998; Ghaffarian and Shahriari, 2017) are a non-solution to the above. Such ML techniques would meet our criteria for configurability and composability; but they would fail the first and second criteria, auditability and specific assumptions that can later be vetted by human developers and human users.
Cargo Sherlock: an SMT-based approach.
@@
To address these goals, we propose Cargo Sherlock, a design and implementation for quantifying supply chain trust based on a formalization in satisfiability modulo theories (SMT). Cargo Sherlock can be used directly by end users to quantify trust when selecting dependencies to add to a project, and its features are motivated by this intended use case; see Figure 1 and Section 2.
@@
We begin with an SMT-based formalization of trust that we call the minimum trust problem (Section 3). This particular formalization assumes that we have a list of assumptions that the user might make about the code (such as: “if the code has a large number of downloads, then it is safe” or “if the code is audited by a trusted organization, then it is safe”); each assumption has an associated cost, which can be thought of as the cost inherent in making that assumption. These assumptions are motivated by prior empirical research; importantly, they are not necessarily true facts (if they were true unconditionally then trust would not be required!), but rather, they are real thought processes that users might potentially employ to conclude that dependencies are safe, and are configurable by users. The minimum trust problem then is to determine the minimum cost of a set of assumptions, encoded in propositional logic, that logically entail that the code is safe. We argue that this model can meet the goals we outlined above; for example, it produces trust costs that are specific to a particular set of assumptions, can be configured to meet specific users’ needs, and are auditable by human developers. We consider two algorithms for solving the minimum trust problem, based on a reduction to SMT. We also show that the problem is NP-hard in the worst case. This formalization is similar in principle to existing work from the logic programming and databases communities, for example, the semiring framework for database provenance (Green and Tannen, 2017), assigning weights to logical conclusions in Datalog (Bistarelli et al., 2008), and solvers for weighted Max-SAT (Heras et al., 2008; Li and Manya, 2009).
@@
For the design and implementation (Section 4), Cargo Sherlock gathers software supply chain metadata and the results of automated program analysis tools to build a first-order relational structure of the software supply chain – that is, of all the dependencies of a given package. It then instantiates the minimum trust problem above on a set of possible assumptions on the variables involved in this first-order structure. The implementation also supports “parameterized assumptions”, where the cost depends in a continuous way on a particular piece of metadata, and “negative assumptions”, which can be used to incorporate information about software components that are distrusted, such as known vulnerabilities or bugs found by the program analysis tools.
@@
Our evaluation (Section 5) asks four questions: Can Cargo Sherlock be used to prevent synthetically generated supply chain attacks? Would Cargo Sherlock identify risky Rust dependencies (crates) that have been involved in known supply chain incidents? How does Cargo Sherlock fare on known vulnerable crates? Lastly, does the performance of Cargo Sherlock scale to Rust crates with many dependencies? Among our results, we find that synthetically generated typosquatting attacks generally result in much higher trust costs. We also find that of the two algorithms we propose, our more optimized algorithm provides timely results for crates with a large number of dependencies.
@@
One way to put these results into context is to think about a particular user model of how we expect users to interact with our tool. For example, due to the higher trust costs, if using our tool in the prescribed model, a user would catch any of the typosquatting attacks considered in our evaluation by identifying and omitting the risky dependency prior to its inclusion.
@@
  • A formal model of the supply chain trust problem and two algorithms for its solution using the SMT theory of linear arithmetic (Section 3);
  • A design and architecture for a usable tool based on these ideas, including concrete input data about Rust code and a collection of built-in candidate assumptions motivated by human factors and available program analyses (Section 4);
  • An evaluation of the efficacy of our tool for identifying synthetic and real-world supply chain attacks, and its scalability on Rust libraries with many dependencies (Section 5).
We conclude the paper with related work (Section 6). Cargo Sherlock is open source (MIT license) and available via a virtual machine image on Figshare [^1] https://doi.org/10.6084/m9.figshare.31094197 and as a repository on GitHub. [^2] https://github.com/davispl/cargo-sherlock
@@ 2. Motivating Overview
Figure 1. Cargo Sherlock example output on the anyhow crate. (Note that lower trust cost, higher distrust cost is better.)
Example output of Cargo Sherlock on the anyhow-1.0.97 crate, showing that it is marked as SAFE due to presence of an audit.
This section presents a motivating overview of Cargo Sherlock from a user-facing perspective; the reader who is more interested in formal details may skip to Section 3. We consider a hypothetical example of Ferris, a Rust developer who wishes to install a Rust crate called aho_corasick (implementing the Aho-Corasick algorithm), and explore how Cargo Sherlock’s approach might help Ferris attain the goals laid out in the introduction.
@@ 2.1. Automating trust costs
After finding the crate aho_corasick from a Google search, Ferris faces a choice: do they trust the crate blindly, or do they first try to determine if the crate is safe to install on their system? If Ferris is a typical developer, they might employ a common middle ground by relying on various heuristics to argue that the crate is likely safe: e.g., that the crate has many downloads, or that it is developed by someone who is trustworthy. At its core, our tool supports Ferris by attempting to automate this reasoning. The premises of Ferris’ arguments are based on assumptions about the supply chain ecosystem (e.g., “the more downloads a crate has, the safer it is”); Cargo Sherlock includes a default list of such assumptions motivated by prior research (see Section 4.3). Ferris may find some assumptions more reasonable than others; we mark each assumption with a cost to reflect its reasonableness. If Ferris runs Cargo Sherlock on aho_corasick, it will present them with aho_corasick’s trust cost: the total cost of the argument whose premises have the lowest total cost to assume (informally, the strongest argument). After using Cargo Sherlock, Ferris determines that aho_corasick is marked as trusted, and that it was marked SAFE due to having a high number of downloads. See Figure 1 for an example output; the trust cost (lower is better) is shown at the top in green.
@@ 2.2. Trust should be specific
Now suppose that Ferris becomes concerned about trusting crates with a high number of downloads. To address this concern, Ferris hires an expert auditor, Crabby, who combs through every line of source code of aho_corasick. Since Ferris trusts Crabby’s expertise, Ferris assigns a very low cost to the assumption that “if the crate is audited by Crabby, it is safe” – lower than any other available assumption. Next, suppose Crabby comes back after the audit and reports that the crate is completely safe to use. In this scenario, it would seem that the previous assumptions used to mark the crate trusted – such as its high number of downloads or the author of the crate – are no longer relevant to Ferris. Since Cargo Sherlock only reports the minimum trust cost for a crate with the associated assumptions, it does not present other (potentially unhelpful) assumptions. In other words, Ferris is only presented with a single specific reason to trust the crate. In Figure 1, the specific assumptions are shown under “Assumptions for Trusting” in green.
@@ 2.3. Trust should be auditable and configurable
Unfortunately, Ferris later finds out that Crabby is untrustworthy. Since this renders the existing audit suspect, Ferris revisits the previous assumptions, and increases the trust cost for “if the crate is audited by Crabby, it is safe”. In turn, Cargo Sherlock will update its result: the trust cost for aho_corasick will increase, and a new minimum set of assumptions will be identified. This same mechanism can be used if Ferris wants to add new information to the tool (e.g., adding additional trusted authors and auditors).
@@ 2.4. Trust should be composable
During this process, Ferris discovers that aho_corasick depends on a second crate, memchr, that was previously overlooked. Since this could present an additional attack surface, Ferris believes that the trust cost for aho_corasick should be dependent on the cost of its dependencies in some cases. One way to model this is with assumptions such as the following one: “if a crate’s dependencies are all safe to use, and the crate does not have any code flagged by a static analysis tool, then the crate is safe.” Cargo Sherlock uses a similar assumption to allow arguments that compose trust costs on a crate’s dependencies to obtain a trust cost for the crate itself.
@@ 2.5. Incorporating known risks
Finally, Ferris discovers the RustSec Advisory Database (Rust Secure Code Working Group, 2025), which lists vulnerabilities for different Rust crates. Ferris wants to check whether aho_corasick is present in the database. Unfortunately, nothing we have considered so far accounts for negative factors that would decrease trust in a crate; intuitively, this is because we have only incorporated assumptions that are about trustworthiness. We thus extend the tool with a distrust model to quantify known risks; we combine the results of the trust and distrust models to obtain a final label for the crate. In Figure 1, the distrust model is shown alongside the trust model with output in red.
@@ 3. Minimum Trust Problem
In this section, we present the formal model behind Cargo Sherlock, which aims to satisfy the requirements from Sections 1 and 2. We define the minimum trust problem formally: given a set of candidate assumptions and a conclusion, it asks to find the set of assumptions with minimum cost that entails the conclusion. We begin with an example, followed by two algorithms for the problem: first, a naive algorithm (Algorithm 1) and second, one based on Horn clause inference (Algorithm 2) that forms the basis for our trust solver implementation. To conclude the section, we show that the complexity of the minimum trust problem for Horn clauses is NP-complete, justifying such SMT-based approaches.